OpenStudy (anonymous):

The ratio of men to woman in a certain factory is 3 to 4. There are 210 men. How many workers are there?

OpenStudy (anonymous):

490

OpenStudy (anonymous):

is it correct....

OpenStudy (anonymous):

M/W = 3/4 W = (M*4)/3 W = (210*4)/3 = 280 Total Workers = M + W = 210 + 280 = 490